Disposal of Fixed Asset

Equipment acquired on January 6 at a cost of $441,600 has an estimated useful life of 10 years and an estimated residual value of $57,600.

a. What was the annual amount of depreciation for the Years 1-3 using the straight-line method of depreciation?

Year Depreciation Expense
Year 1 $
Year 2 $
Year 3 $

b. What was the book value of the equipment on January 1 of Year 4?
$

Feedback

c. Assuming that the equipment was sold on January 3 of Year 4 for $310,100, journalize the entry to record the sale. If an amount box does not require an entry, leave it blank.

Jan. 3 Cash
Accumulated Depreciation-Equipment
Loss on Sale of Equipment
Equipment

Feedback

d. Assuming that the equipment had been sold on January 3 of Year 4 for $332,900 instead of $310,100, journalize the entry to record the sale. If an amount box does not require an entry, leave it blank.

Jan. 3 Cash
Accumulated Depreciation-Equipment
Equipment
Gain on Sale of Equipment
Answer #1
a
Year Depreciation Expense
Year 1 38400 =(441600-57600)/10
Year 2 38400 =(441600-57600)/10
Year 3 38400 =(441600-57600)/10
b
Book value of the equipment 326400 =441600-(38400*3)
c
Cash 310100
Accumulated Depreciation-Equipment 115200 =38400*3
Loss on Sale of Equipment 16300
         Equipment 441600
d
Cash 332900
Accumulated Depreciation-Equipment 115200
         Equipment 441600
         Gain on Sale of Equipment 6500